OpenStudy (anonymous):

]Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (anonymous):

Do you know point-slope formula?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

It's y-y(initial)=m(x-x(initial))

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

m=slope

OpenStudy (anonymous):

x initial is the x value you were given. y intial is the y value you were given.

OpenStudy (anonymous):

SO, y+1=4(x+3)